### v1

# Mechanisms vs. Outcomes: Probing for Syntax Fails to Explain Performance on Targeted Syntactic Evaluations

Abstract
Large Language Models (LLMs) exhibit a robust mastery of syntax when processing and generating text. While this suggests internalized understanding of hierarchical syntax and dependency relations, the precise mechanism by which they represent syntactic structure is an open area within interpretability research. Probing provides one way to identify the mechanism of syntax being linearly encoded in activations, however, no comprehensive study has yet established whether a model’s probing accuracy reliably predicts its downstream syntactic performance. Adopting a “mechanisms vs. outcomes” framework, we evaluate 32 open-weight transformer models and find that syntactic features extracted via probing fail to predict outcomes of targeted syntax evaluations across English linguistic phenomena. Our results highlight a substantial disconnect between latent syntactic representations found via probing and observable syntactic behaviors in downstream tasks.
Mechanisms vs. Outcomes: Probing for Syntax Fails to Explain Performance on Targeted Syntactic Evaluations
Ananth Agarwal, Jasper Jian, Christopher D. Manning, Shikhar Murty Stanford University {ananthag, manning, smurty}@cs.stanford.edu, jjian@stanford.edu
1 Introduction
The remarkable ability of LLMs to process and generate text that respects a rich diversity of syntactic constraints strongly suggests that models have sophisticated syntactic knowledge. Yet our understanding of internal representations of these syntactic facts is lacking, which has motivated research into interpretability paradigms searching for syntactic mechanisms, such as mechanistic interpretability and causal abstractions Geiger et al. (2022); Murty et al. (2022). We study probing, a prominent approach in this domain which assumes that model activations are the mechanism through which syntactic knowledge is encoded and that this knowledge is linearly recoverable using small supervised models called probes (Conneau et al., 2018; Hewitt and Manning, 2019; Tenney et al., 2019, among others). However, no comprehensive study has evaluated whether probing accuracy is indicative of targeted syntactic outcomes.
Figure 1: Mechanisms vs. outcomes setup. We find no convincing predictive power of syntax probing accuracy on downstream syntactic evaluation accuracy. (a) Mechanism: probe $g_{\phi}$ extracts a dependency parse tree from $S_{acc}$ word-level hidden states $h_{i}$; UUAS = 4/6 edges in this toy example. (b) Outcome: evaluate minimal pair accuracy.
We deepen understanding of the relationship between syntactic mechanisms and outcomes by asking: Do probing accuracies, measured by tree attachment scores of syntactic probes, effectively predict a model’s downstream performance on targeted syntactic evaluations? We evaluate syntactic outcomes using BLiMP (Warstadt et al., 2020), a benchmark of minimal pairs spanning well-established English grammatical contrasts. Assigning higher probability to the acceptable sentence in an acceptable-unacceptable pair has been established as a desired linguistic outcome Marvin and Linzen (2018); Warstadt et al. (2020).
We train three syntax probes on 32 open-weight transformer models. We then fit Ordinary Least Squares (OLS) regressions modeling BLiMP minimal pairs accuracy as a function of probe attachment scores on the grammatically acceptable sentences (Fig 1). We demonstrate that across all levels of analysis—amalgamation of all BLiMP linguistic phenomena, individual phenomena, and targeted suites of challenging minimal pairs—syntax probing accuracy shows no clear predictive power for downstream outcomes. [^1] Code will be uploaded to GitHub soon.
Concretely, we find that probing performance—measured via directed and undirected unlabeled attachment scores (UAS/UUAS)—fails to predict model accuracy on BLiMP minimal pairs. At the full-dataset level, regression fits are negligible, with no statistically significant relationships across 32 models and three probes. At the per-phenomenon level, only anaphor agreement shows a significant correlation, but follow-up controls reveal that even a non-syntactic probe can produce similar effects, suggesting the result is likely spurious. In finer-grained evaluations targeting challenging syntactic paradigms, we test whether probes can recover critical edges—such as the subject-to-verb dependency in agreement examples—that are plausibly relevant for resolving minimal pairs. However, even when the probe identifies these edges correctly, model success on the corresponding minimal pair is only weakly aligned (40–60% match rate across models). Finally, at the sentence level, UUAS scores are largely overlapping between correctly and incorrectly evaluated minimal pairs in over 85% of BLiMP suites. Taken together, our results suggest that while syntactic probing is popular, it fails to robustly reveal a model’s latent syntactic knowledge—external targeted evaluations remain the gold standard.
2 Setup
Given an input sentence $S\triangleq(w_{1},\ldots,w_{N})$, a pre-trained transformer $f_{\theta}$ produces contextual vectors $h_{i}\in\mathbb{R}^{d{}}$ for each word $w_{i}$ at every layer. Our probes $g_{\phi}{}:\mathbb{R}^{d{}}\to\mathbb{R}^{k{}}$, $k{}\leq d{}$ are learned functions that project contextual hidden states linearly into a $k{}$-dimensional sub-space. Linear and bilinear probes have been shown to achieve high selectivity; they effectively capture linguistic properties of the probed representation (Hewitt and Liang, 2019).
2.1 Mechanisms
We train three syntax probes with different optimization objectives for comprehensive coverage. We evaluate the correctness with which each probe extracts the dependency parse of a sentence.
Structural Probe.
Proposed in Hewitt and Manning (2019), $g^{\text{struct}}_{\phi}$ is a linear transformation ($\phi=\{B^{\text{struct}}\in\mathbb{R}^{d{}\times k{}}\}$) that learns to encode the tree distance $d_{ij}$ between word pairs $w_{i}$ and $w_{j}$ in the probe’s projected space by minimizing
$g^{\text{struct}}_{\phi}(h)=B^{\text{struct}}h$
$\min_{\phi}\sum_{S}\frac{1}{|S|^{2}}\sum_{i,j}\left|d_{ij}-\|g^{\text{struct}}_{\phi}(h_{i}-h_{j})\|_{2}^{2}\right|.$ (1)
The test metric is undirected unlabeled attachment score (UUAS): the fraction of gold tree edges included in the minimum spanning tree calculated from the probe’s predicted tree distances. Punctuation and the root relation are excluded.
Orthogonal Structural Probe
Limisiewicz and Mareček (2021) replace the $g^{\text{struct}}_{\phi}$ linear transformation with an orthogonal transformation and scaling vector $g^{\text{ortho}}_{\phi}$ where parameters of the probe $\phi=\{V\in\mathbb{R}^{d{}\times d},\bar{d}\in\mathbb{R}^{d}\}$. $V$ is orthogonal, and the authors maintain its orthogonality during training using Double Soft Orthogonality Regularization (DSO) Bansal et al. (2018) with regularization $\lambda_{o}=0.05$. The overall probe training objective is
$g^{\text{ortho}}_{\phi}(h)=\bar{d}\odot Vh$
$\min_{\phi}\sum_{S}\frac{1}{|S|^{2}}\sum_{i,j}\left|d_{ij}-\|g^{\text{ortho}}_{\phi}(h_{i}-h_{j})\|_{2}^{2}\right|$
$+\lambda_{o}\text{DSO}(V).$ (2)
Head Word Probe
Inspired by Clark et al. (2019), $g^{\text{head}}_{\phi}$ ($\phi=\{B^{\text{head}}\in\mathbb{R}^{d{}\times k{}}\}$) formulates a classification task with cross-entropy loss: for each $w_{i}$, including punctuation, predict the head word in $\{\text{ROOT},w_{1},\ldots,w_{N}\}\setminus\{w_{i}\}$. Let $H$ be cross entropy loss and $\text{head}_{i}$ be the head word for $w_{i}$:
$g^{\text{head}}_{\phi}(h)=B^{\text{head}}h$
$\hat{d}_{ij}=\|g^{\text{head}}_{\phi}(h_{i}-h_{j})\|_{2}$
$\min_{\phi}\sum_{S}\frac{1}{|S|}\sum_{i}H(\{\hat{d}_{ij}\mid i\neq j\},\text{head}_{i}).$ (3)
Since the edge prediction is directed, the test metric is unlabeled attachment score (UAS).
Control Probe.
Importantly, training syntactic probes alone is insufficient to draw conclusions about their ability to predict syntactic evaluation performance because the quality of model hidden states is a confounding variable affecting both probe and syntactic evaluation performance. To isolate this confounder, we propose a simple control probing task where performance depends only on hidden state quality and not on latent syntactic structure. Concretely, for each sentence, we form all word pairs where the treebank-specific part-of-speech (XPOS) of the two words is the same and train a probe to recover the $L_{2}$ distance between the two words’ GloVe vectors Pennington et al. (2014). Let $v_{i}$ denote the GloVe vector of $w_{i}$. The training objective for the control probe is
$\tilde{d}_{ij}=\|g^{\text{ctrl}}_{\phi}(h_{i}-h_{j})\|_{2}$
$\min_{\phi}\sum_{S}\frac{1}{|S|}\sum_{\mathclap{\begin{subarray}{c}(i,j)\\ \text{POS}(w_{i})=\text{POS}(w_{j})\end{subarray}}}\;\text{Huber}(\tilde{d}_{ij},\|v_{i}-v_{j}\|_{2}).$ (4)
We use Huber loss for outlier robustness. The evaluation metric for the probe is the Spearman correlation $\rho_{s}$ between predicted and actual GloVe distances.
This probe is trained to recover non-syntactic information as modeled by an uncontextualized word embedding model, GLoVe, whose vector geometries have been shown to encode linguistic information like lexical semantics (Brown et al., 2023). We further ensure that our control probe is not indirectly capturing syntactic part-of-speech information by only training the probe to recover within word-pair distances where the words share the same XPOS. Crucially, any predictive success of the control probe on syntactic evaluations warrants a more cautious interpretation of the results of syntactic probes, as it suggests that the latter’s explanatory power may not be attributable solely to the recoverability of syntactic representations.
2.2 Outcomes
Our benchmark for linguistic knowledge outcomes is BLiMP accuracy. BLiMP has 67 template-generated sub-datasets called paradigms grouped into 13 linguistic phenomena across English morphology, syntax, and semantics Warstadt et al. (2020). [^2] The two semantic phenomena, quantifiers and NPI licensing, still require knowledge of hierarchical syntactic structure, e.g., to determine scope for NPI licensing (Ladusaw, 1979). Each paradigm has 1000 acceptable-unacceptable minimal pairs $(S_{acc},S_{unacc})$. Accuracy on a set of pairs $\mathcal{D}$ is computed as
$\frac{1}{|\mathcal{D}|}\sum_{\mathcal{D}}\mathbb{1}\{\log P_{f_{\theta}}(S_{acc})>\log P_{f_{\theta}}(S_{unacc})\}.$ (5)
For decoder models, $\log P_{f_{\theta}}(S)$ is the sum of token level log-probabilities. For encoder and encoder-decoder models, we instead use pseudo-log-likelihood scores (PLL; Salazar et al., 2020).
2.3 Models
We test 32 open-weight pretrained decoder, encoder, and encoder-decoder models up to 8B parameters. Appendix A has the full list.
3 Experimental Setup
Probe Training.
We train probes on Stanza Qi et al. (2020) Universal Dependencies Zeman et al. (2019) dependency parses of the Penn Treebank (PTB; Marcus et al., 1993) using its standard train and validation splits. We report training parameters and compute in Appendix B. $g^{\text{struct}}_{\phi}$ and $g^{\text{head}}_{\phi}$ have fixed probe dimension $k{}=256$, while $g^{\text{ortho}}_{\phi}$ has $k=d$ for each model to ensure orthogonality is feasible. Let $L$ be the number of layers in LM $f_{\theta}$. To reduce the influence of a word’s semantic properties, we subtract word embeddings from each layer’s contextual states
$h_{i}^{\ell}=h_{i}^{\ell}-h_{i}^{0}\hskip 14.22636pt\forall\ell\in\{1,\ldots,L\}.$
Training produces $L$ different probes for each model. For each of $g^{\text{struct}}_{\phi}$ and $g^{\text{head}}_{\phi}$, we select the probe with the best PTB test metric. For $g^{\text{ortho}}_{\phi}$, given limited compute resources and the fact that it is derived from $g^{\text{struct}}_{\phi}$, we train a single probe per model on the layer with the best $g^{\text{struct}}_{\phi}$ test metric. For $g^{\text{ctrl}}_{\phi}$ we train at most two probes per model corresponding to the best layers for $g^{\text{struct}}_{\phi}$ and $g^{\text{head}}_{\phi}$, which may match.
Control Probe Validation.
We construct a simple evaluation dataset to test how well our control probe abstracts away contextual linguistic information to verify its design. For 500 random words chosen from SimLex999 (Hill et al., 2014), we prompt GPT-4o OpenAI et al. (2024) to generate 10 sentences that contain the word in substantially different contexts. Refer to Appendix C for the prompt and sample sentences.
We first measure the variance of each word’s contextual hidden states (at the layer $g^{\text{ctrl}}_{\phi}$ is trained at) across the 10 samples, and report the average across all 500 words. We then apply $g^{\text{ctrl}}_{\phi}$ on these hidden states and recompute this quantity. If the probe successfully abstracts away contextually-determined linguistic information like syntax and retains non-contextual information like lexical semantics, we expect the mean variance of a given word hidden state to significantly decrease, since lexical semantics of a given word should remain the same regardless of context.
Mechanisms vs. Outcomes Regression.
Our objective is model-level mechanisms vs. outcomes correlation because we want to study how effectively the method of syntactic probing reveals an arbitrary model’s latent syntactic knowledge. We use the PTB test set scores to choose the best probe from each probe family and apply it to every $S_{acc}$ from BLiMP to compute UUAS/UAS/$\rho_{s}$ as predictor variables (mechanisms). $S_{acc}$ and $S_{unacc}$ feed into minimal pair evaluation as the response variable (outcomes).
We fit simple (eqn 6) and multiple (eqn 7) OLS regressions.
$y=\beta_{0}+\beta_{1}x_{1}+\epsilon$ (6)
$y=\beta_{0}+\beta_{1}x_{1}+\beta_{2}x_{2}+\epsilon$ (7)
$y$ is minimal pairs accuracy (eqn 5). We run simple regression with each probe individually: $x_{1}$ is either BLiMP UUAS, UAS, or Spearman correlation $\rho_{s}$. For multiple regression, $x_{1}$ is UUAS or UAS and $x_{2}$ is $\rho_{s}$. Comparing $\beta_{1}$ and its $p$-value for the syntax probe when it is the sole predictor versus when paired with the control reveals the extent to which it predicts grammaticality outcomes. Since the two regression models are nested, we also include log-likelihood ratio tests (LRT) for robust statistical comparisons.
We experiment at three BLiMP granularities: (1) full dataset (average results of all test suites), (2) each of the 13 phenomena separately with Holm-Bonferroni correction, (3) the two test suites in each of the subject-verb agreement and filler-gap phenomena that scored lowest in Warstadt et al. (2020). For the individual test suite experiments, rather than tree UUAS/UAS, we determine if the induced syntactic tree contains a critical edge corresponding to the phenomenon targeted by the minimal pairs. Appendix D discusses critical edge criteria. We hypothesize that the binary outcome of the probe’s ability to extract the critical edge on $S_{acc}$ is linked to the minimal pair binary outcome.
4 Results & Discussion
Attachment score typically peaks in the middle layers.
Fig 2 plots PTB test $g^{\text{struct}}_{\phi}$ UUAS across model layers. Best layers for GPT-2 (Radford et al., 2019) match with previous results from the literature (Eisape et al., 2022), validating our probe training procedure. In Appendix E, Fig 13 plots PTB test $g^{\text{head}}_{\phi}$ UAS, and we further show that our syntax probes’ abilities to recover encoded parse tree information in BLiMP sentences correlate well across training objective ($g^{\text{struct}}_{\phi}$ vs. $g^{\text{head}}_{\phi}$, Fig 14) and architecture ($g^{\text{struct}}_{\phi}$ vs. $g^{\text{ortho}}_{\phi}$, Fig 15). Furthermore, the comparable range of UUAS/UAS scores between the PTB test set and BLiMP shows the probes generalize from human-written PTB text to template-generated BLiMP.
Control probe erases contextual information.
Fig 3 shows that for all models, regardless of the scale of the variance in the average word’s contextual hidden states across sentences, the variance in the control probe’s projected representation space is near-zero. This shows the probe nullifies context and uncovers GloVe-like linguistic information within hidden states.
Figure 2: Penn Treebank test set structural probe UUAS for each layer of a sample of our models. The star icon for a model indicates the layer with the best test set accuracy that is used for BLiMP evaluation. For most models, this occurs in the first half. Our results for GPT-2 117M (small) and GPT-2 1.6B (XL) closely match those of Eisape et al. (2022), supporting the correctness of our implementation. For T5 3B, although most layers yield low UUAS, the best layer (13) still exceeds 0.7.
Figure 3: Control probes consistently erase contextual information in hidden states, as evidenced by near-zero variance of the previously-contextual word contextual hidden states in the projected representation space.
Figure 4: Simple regression plots for $g^{\text{struct}}_{\phi}$ (top row), $g^{\text{ortho}}_{\phi}$ (middle row), and $g^{\text{head}}_{\phi}$ (bottom row). Each panel is annotated with adjusted $R^{2}$ and the $p$-value of $\beta_{1}$. Per-phenomenon results have Holm-Bonferroni correction. The first column of panels shows that at the full dataset granularity, no probe explains the spread in minimal pairs accuracy with any statistical significance. At per-phenomenon granularity, the second column contains the phenomenon with the lowest $p$-value per probe. We additionally highlight subject-verb agreement (third column) and filler-gap (fourth column) as strongly syntactic tasks with critical edges that we identify in Appendix D. Refer to Appendix F for complete simple regression results.
Dataset Simple Regression ($x_{1}=g^{\text{struct}}_{\phi}$) Multiple Regression ($x_{1},x_{2}=g^{\text{ctrl}}_{\phi}$) LRT
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
Overall 0.133 0.117 0.049 0.109 0.214 0.051 1.161 0.281
Anaphor Agreement 0.165 0.019 0.268 0.168 0.040 0.243 0.027 1.0
Subject Verb Agr -0.108 1.0 0.007 -0.047 1.0 0.128 5.236 0.243
Filler Gap Dependency -0.028 1.0 -0.033 0.033 1.0 -0.05 0.546 1.0
Table 1: Comparison of simple and multiple regression statistics with $x_{1}=$ $g^{\text{struct}}_{\phi}$ UUAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon $p$-values are Holm-Bonferroni corrected and bold text indicates statistical significance ($p$ < 0.05). While the simple regression model is a better fit for anaphor agreement than the multiple regression model, there is no clear signal across any other phenomena. Refer to Table 5 in Appendix F for complete simple-multiple regression statistics for all phenomena.
Dataset Simple Regression ($x_{1}=g^{\text{ortho}}_{\phi}$) Multiple Regression ($x_{1},x_{2}=g^{\text{ctrl}}_{\phi}$) LRT
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
Overall 0.127 0.157 0.035 0.102 0.266 0.041 1.299 0.254
Anaphor Agreement 0.173 0.002 0.364 0.173 0.005 0.342 0.004 1.0
Subject Verb Agr -0.119 1.0 0.014 -0.059 1.0 0.132 5.183 0.251
Filler Gap Dependency -0.1 1.0 -0.027 -0.058 1.0 -0.049 0.402 1.0
Table 2: Comparison of simple and multiple regression statistics with $x_{1}=$ $g^{\text{ortho}}_{\phi}$ UUAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon $p$-values are Holm-Bonferroni corrected and bold text indicates statistical significance ($p$ < 0.05). While the simple regression model is a better fit for anaphor agreement than the multiple regression model, there is no clear signal across any other phenomena. The higher capacity of $g^{\text{ortho}}_{\phi}$ relative to $g^{\text{struct}}_{\phi}$ does not translate to improved regression fit, suggesting that its additional probe dimensions are extra. Refer to Table 6 in Appendix F for complete simple-multiple regression statistics for all phenomena.
Dataset Simple Regression ($x_{1}=g^{\text{head}}_{\phi}$) Multiple Regression ($x_{1},x_{2}=g^{\text{ctrl}}_{\phi}$) LRT
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
Overall 0.0 0.998 -0.033 -0.048 0.700 0.091 5.194 0.023
Binding 0.534 0.064 0.209 0.522 0.108 0.184 0.085 1.0
Subject Verb Agr -0.205 1.0 0.041 -0.221 1.0 0.025 0.56 1.0
Filler Gap Dependency -0.125 1.0 -0.028 -0.15 1.0 -0.055 0.256 1.0
Table 3: Comparison of simple and multiple regression statistics with $x_{1}=$ $g^{\text{head}}_{\phi}$ UAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon $p$-values are Holm-Bonferroni corrected and bold text indicates statistical significance ($p$ < 0.05). The LRT in the first row indicates the superiority of the multiple regression model, but its adjusted $R^{2}$ is still very weak. Refer to Table 7 in Appendix F for complete simple-multiple regression statistics for all phenomena.
No significant fit on full BLiMP.
The first column of Fig 4 shows OLS simple regression result plots at the full BLiMP dataset granularity. Each of the three syntax probes demonstrates a minimal effect size with none reaching statistical significance. $g^{\text{head}}_{\phi}$ has no measurable effect at all. Its UAS values generally exceed the UUAS values of other probes, reflecting successful learning of relation directionality, but there is insufficient stratification along the y-axis for any predictive power.
Table 1 ($g^{\text{struct}}_{\phi}$), Table 2 ($g^{\text{ortho}}_{\phi}$), and Table 3 ($g^{\text{head}}_{\phi}$) compare simple and multiple regression statistics. For $g^{\text{head}}_{\phi}$, the significant $p$-value of the LRT conveys that considering non-syntactic signal from activations provides better fit on full BLiMP than relation direction-oriented signal alone, but the adjusted $R^{2}$ is still very weak.
Only anaphor agreement has statistical significance at the per-phenomenon granularity.
The second, third, and fourth columns of panels in Fig 4 show simple regression result plots for anaphor agreement, subject-verb agreement, and filler-gap dependencies respectively. Out of the three syntax probes and all phenomena (full regression tables in Appendix F), the only statistically significant correlations are $g^{\text{struct}}_{\phi}$ and $g^{\text{ortho}}_{\phi}$ on anaphor agreement. Probes show poor fit on minimal pairs accuracy for all other phenomena. Strikingly, probes show no predictive power even for phenomena like subject-verb agreement and filler-gap dependencies whose behavior we might expect to be closely reliant on syntactic representations.
However, the anaphor agreement result warrants caution. Fig 5 shows the surprising result that the control probe, which is trained on a non-morphosyntactic task, can be highly correlated to the BLiMP irregular forms morphosyntactic task. Non-syntactic probes can predict syntactic benchmark performance. In the absence of broad systematic correlations, one-off probe/task correlations are not evidence of syntactic representations in hidden states being a primary explanatory variable for downstream behavior, even at the per-phenomenon level: probes designed to be unrelated can be correlated to single tasks. Even with Holm-Bonferroni correction, the anaphor agreement fit could be an artifact of multiple statistical testing. As for the full BLiMP results, Table 1 ($g^{\text{struct}}_{\phi}$), Table 2 ($g^{\text{ortho}}_{\phi}$), and Table 3 ($g^{\text{head}}_{\phi}$) present the simple and multiple regression statistics side-by-side. Aside from anaphor agreement, there is no consistent result of the simpler regression model providing a statistically significant better fit than the regression that includes the control. As a further example, $g^{\text{head}}_{\phi}$ encodes hierarchical dominance of the head word, which we would expect to influence NPI licensing; however, Table 7 shows no such fit.
Figure 5: Control $g^{\text{ctrl}}_{\phi}$ unexpectedly achieves statistical significance for predicting the BLiMP irregular forms phenomena, which is a morphosyntactic task.
Moderate mismatch rate in finer-grained test suite evaluation.
Fig 6 shows that for 3 out of the 4 challenging subject-verb agreement and filler-gap paradigms discussed in Section 3, the match rate between $g^{\text{struct}}_{\phi}$ correctly predicting the critical edge and the model resolving the minimal pair correctly is between 40–60% for the majority of models. The scatter plots show a substantial spread in model performance on both axes, which suggests the difficulty of the tasks. However, high Hamming distances between correct prediction binary outcomes mean that errors do not align as much as expected. We conclude that extracting these critical edges, despite its possible relevance to deciding between the words which are different between the minimal pairs (Appendix D), does not in fact predict evaluation success on the minimal pair. Fig 16 yields the same conclusion from $g^{\text{head}}_{\phi}$ probe results.
Figure 6: Challenging subject-verb agreement and filler-gap phenomena critical edge prediction results for $g^{\text{struct}}_{\phi}$. Hamming distance is computed between the binary outcome of the structural probe predicting the critical edge correctly in $S_{acc}$ and the binary outcome of the model resolving the minimal pair correctly. Successful critical edge prediction does not translate to successful minimal pair resolution for 3 out of 4 test suites. The majority of models show mismatches occurring in approximately 40% to 60% of cases (Hamming distance buckets 0.4 to 0.6). Generally, the harder the test suite, indicated by clustering of points lower on the y-axis in the scatter plots, the higher the Hamming distance.
UUAS score distributions largely overlap for correct vs. incorrect minimal pairs.
While cross-model correlations between probing and downstream performance have proved elusive, a natural follow-up investigation for each model is the sentence-level comparison of if there is a statistically significant difference in the UUAS score distribution between the pool of sentences where the minimal pair is evaluated correctly and the pool where it is evaluated incorrectly. For each BLiMP test suite, we run a two-sample $t$-test for each model on these pools of UUAS scores and apply Holm-Bonferroni correction on the test $p$-values. The results in Fig 7 show that only 8 out of 67 BLiMP test suites have statistical evidence for the mean of UUAS scores for correct minimal pairs being higher than that of incorrect minimal pairs for at least 5 models. Yet again, we conclude that for most models and test suites, UUAS score of the probe at the sentence level (mechanism) does not predict if the model will succeed at the corresponding minimal pair evaluation (outcome).
Figure 7: Sentence-level UUAS score distribution $t$-test results. Out of 67 BLiMP test suites, only the 26 shown here have at least one model that has a statistically significant difference in the UUAS score means. The $t$-test provides statistical evidence that the mean of UUAS scores where the minimal pair is evaluated correctly is greater than the mean where it is evaluated incorrectly.
5 Related Work
Our work builds on the extensive literature examining the syntactic capabilities of language models through probing (Hewitt and Manning, 2019; Clark et al., 2019; Müller-Eberstein et al., 2022; Limisiewicz and Mareček, 2021; Diego-Simón et al., 2024, among others).
Tucker et al. (2021) demonstrate that counterfactual interventions on contextual embeddings—designed to erase information identified by a probe—can affect downstream behavior in predictable ways. However, they also find that such effects are not robust across models or probes. Similarly, Eisape et al. (2022) perform causal intervention on transformer states during autoregressive generation from GPT-2. Here, we conduct a more comprehensive study to understand the connection between probing and downstream syntactic behavior across model types, scale, and probe categories. Ruzzetti et al. (2023) take a different approach from probing to explore the linguistic properties of monolingual BERT Devlin et al. (2019). While our work is currently limited to English, Ruzzetti et al. (2023) compare layer weight matrices across monolingual models to uncover linguistically-motivated typological similarities. Their work can serve as a blueprint for future multilingual probing evaluations since common phenomena across languages may be predictable from probes trained on monolingual models. Prior work (Bommasani et al., 2022; Belinkov, 2022) cautions against inferring causality between probed representations and downstream tasks Ravichander et al. (2021); Elazar et al. (2021); Belinkov (2022). Belinkov (2022) surveys probing research, emphasizing the importance of careful task and control design for meaningful insights.
6 Conclusion
Targeted evaluations like BLiMP minimal pairs reveal the extent of model syntactic knowledge. High BLiMP scores observed across the open-weight models studied here suggest effective learning of syntactic phenomena. Our work shows that the mechanism of recovering syntactic knowledge from model activations through linear probing does not at all predict targeted evaluation performance. Despite extensive research on probing, further work is necessary to develop a robust understanding of how syntax is represented within models.
7 Limitations
Our evaluation only uses English grammatical benchmarks and could usefully be extended to other languages, but a failure of probe predictive power on English suggests the same result would hold for other languages. Moreover, while we design our syntax probes to be lightweight and to learn to extract syntax with different training objectives, they do not encompass all possible syntactic probing tasks or architectural variants.
Recent probing work aims to make causal claims between probed representations and observed model downstream behavior Belinkov (2022). By intervening to modify these representations, such work investigates if these representations are actually “used” by the model for specific tasks. Our work does not perform causal interventions; instead our focus is the viability of the mechanism conveyed by probing – syntactic knowledge can be linearly decoded from contextual representations – at several experimental granularities. Since our results are primarily negative, there is no discovered correlation to investigate deeper for causality anyway.
Acknowledgements
The authors would like to thank the anonymous reviewers for constructive feedback during the review period. CM is a fellow in the CIFAR Learning in Machines and Brains program. We thank members of the Stanford NLP Group for their insightful comments and support throughout the project.
Appendix A Models
Model HuggingFace ID
Architecture: Decoder
GPT-2 117M Radford et al. (2019) gpt2
GPT-2 345M gpt2-medium
GPT-2 774M gpt2-large
GPT-2 1.6B gpt2-xl
Pythia 410M Biderman et al. (2023) EleutherAI/pythia-410m
Pythia 1.4B EleutherAI/pythia-1.4b
Pythia 2.8B EleutherAI/pythia-2.8b
Pythia 6.9B EleutherAI/pythia-6.9b
GPT-J 6B Wang and Komatsuzaki (2021) EleutherAI/gpt-j-6b
Falcon 7B Almazrouei et al. (2023) tiiuae/falcon-7b
Llama 3.2 1B Grattafiori et al. (2024) meta-llama/Llama-3.2-1B
Llama 3.2 3B meta-llama/Llama-3.2-3B
Llama 2 7B Touvron et al. (2023) meta-llama/Llama-2-7b-hf
Llama 3 8B meta-llama/Meta-Llama-3-8B
Llama 3.1 8B meta-llama/Meta-Llama-3.1-8B
Mistral 7B Jiang et al. (2023) mistralai/Mistral-7B-v0.1
Gemma 2B Team et al. (2024a) google/gemma-2b
Gemma 2 2B Team et al. (2024b) google/gemma-2-2b
OLMo 1B Groeneveld et al. (2024) allenai/OLMo-1B-hf
OLMo 7B Groeneveld et al. (2024) allenai/OLMo-1.7-7B-hf
Phi 2 2.7B Gunasekar et al. (2023) microsoft/phi-2
Architecture: Encoder
ELECTRA-Small 14M Clark et al. (2020) google/electra-small-discriminator
ELECTRA-Base 110M google/electra-base-discriminator
ELECTRA-Large 335M google/electra-large-discriminator
RoBERTa 125M Liu et al. (2019) FacebookAI/roberta-base
RoBERTa 355M FacebookAI/roberta-large
DeBERTa v2 900M He et al. (2021) microsoft/deberta-v2-xlarge
DeBERTa v2 1.5B microsoft/deberta-v2-xxlarge
Architecture: Encoder-Decoder
T5 35M Raffel et al. (2020) google-t5/t5-small
T5 110M google-t5/t5-base
T5 335M google-t5/t5-large
T5 3B google-t5/t5-3b
Table 4: All 32 models used for every experiment.
Appendix B Probe Training Details
All training and inference occurs on a single NVIDIA RTX 6000 Ada Generation or NVIDIA RTX A6000 GPU depending on cluster availability. Training probes on the largest models could take up to 4 days to complete. In the Penn Treebank, the training set has 39,831 sentences, validation set 1,700 sentences, and test set 2,416 sentences. In the following sections, we specify training hyperparameters for each probe.
B.1 Structural Probe ($g^{\text{struct}}_{\phi}$)
  • Batch size: 32
  • Max epochs: 300
  • Early stopping criteria: validation loss; patience of 50 epochs; eval interval of 1 epoch
  • Learning rate: $1e^{-4}$; AdamW optimizer
  • Learning rate schedule: Linear decay with warmup for 10% of max epochs
B.2 Orthogonal Structural Probe ($g^{\text{ortho}}_{\phi}$)
  • Batch size: 32, 64
  • Max epochs: 50
  • Early stopping criteria: validation loss; patience of 5 epochs; eval interval of 1 epoch
  • Learning rate: $1e^{-4}$; AdamW optimizer
B.3 Head Word Probe ($g^{\text{head}}_{\phi}$)
  • Batch size: 32
  • Max epochs: 300
  • Early stopping criteria: validation UAS; patience of 50 epochs; eval interval of 1 epoch
  • Learning rate: $1e^{-4}$; AdamW optimizer
  • Learning rate schedule: Linear decay with warmup for 10% of max epochs
B.4 Control Probe ($g^{\text{ctrl}}_{\phi}$)
  • Batch size: 128
  • Max epochs: 300
  • Early stopping criteria: validation loss; patience of 50 epochs; eval interval of 1 epoch
  • Learning rate: $1e^{-4}$; AdamW optimizer
  • Learning rate schedule: Linear decay with warmup for 10% of max epochs
Appendix C Control Probe Validation
We prompt GPT-4o to elicit sentences with varied contextual usage for a given word:
I will give you a word. Generate 10 sentences that use the word in substantially different contexts. Do not capitalize or add any punctuation to the word, but you can add punctuation elsewhere in the sentences. Do not make the word the first word of the sentence; instead it should be near the end! Return only the list of sentences in the format of a Python list without Markdown. Word: {{word}}
Outputs for the words “window” and “flour”:
"window": [
"The cat loves to sit by the window and watch the birds.",
"After the update, the software has a new window that is more user-friendly.",
"She opened the window to let in some fresh air.",
"There was a brief window of opportunity to complete the task.",
"The display showed a colorful window of stained glass.",
"He peered through the window to see what was happening inside.",
"The startup process will initiate a pop-up window for further instructions.",
"In the story, the hero jumps out of a high window to escape.",
"The interview was scheduled within a tight window of time.",
"She painted a beautiful scene of a sunrise in the window of her art studio."
"flour": [
"To bake the cake, youll need to sift the flour thoroughly.",
"She dusted the countertop with flour before kneading the dough.",
"The recipe calls for almond flour instead of regular wheat.",
"A fine layer of flour covered the bakers apron after a long day.",
"If you want a crispy crust, coat the chicken with flour before frying.",
"They decided to try a gluten-free flour for their cookies this time.",
"The breads texture depends largely on the quality of the flour used.",
"After spilling the bag, a cloud of flour filled the air in the kitchen.",
"He couldnt believe how much more expensive organic flour was at the store.",
"The children made a mess while trying to measure the flour for the pancakes."
Appendix D Critical Tree Edges for BLiMP Phenomena
We investigate if the syntax probe’s accuracy in predicting a critical edge we define on a per-paradigm basis in the acceptable sentence’s parse tree is consistent with the model’s outcome of assigning the acceptable sentence a higher log probability than the unacceptable one. The critical edge involves the minimal difference in the pair’s sentences. In the following sections, we justify the critical edges for the subject-verb agreement and filler-gap paradigms that we experiment with:
  • Subject-verb agreement
    • Distractor agreement relational noun
    • Distractor agreement relative clause
  • Filler-gap dependencies
    • Wh. vs. that with gap
    • Wh. vs. that with gap long distance
D.1 Subject-Verb Agreement
The critical edge is one that connects the relevant subject and the verb targeted by the minimal pair. Consider the following test instance from the distractor agreement relational noun suite:
The prints of every vase aggravate Nina. (acc.)
The prints of every vase aggravates Nina. (unacc.)
Agreement on the verb ‘aggravate’ is determined by its subject, the plural noun ‘prints’. This is represented in parse trees as an edge connecting ‘prints’ and ‘aggravate,’ labeled nsubj. Fig 8 shows the parse tree for the acceptable sentence. If a model’s syntactic representations are related to its success on subject-verb agreement, nsubj is the critical dependency to evaluate – a model successfully representing this edge means that it is minimally aware of what noun is its subject, and thus, which noun it should use to determine whether ‘aggravate’ or ‘aggravates’ should be assigned a higher log probability (see Clark et al., 2019; Jian and Reddy, 2023 for similar discussions of syntactic attention heads).
Under the UD formalism whose parses Stanza produces, the verb which agrees with the subject is not always the head of the nsubj relation in a sentence. For example, in Fig 10 the verb which agrees with the subject is the auxiliary ‘have’, but the UD parse assigns ‘alarmed,’ which does not agree, as the head. For these critical edge experiments only, we filter out all incongruent cases where UD assigns no edge between the subject and critical verb, which leaves 236 valid minimal pairs for distractor agreement relational noun and 345 for distractor agreement relative clause.
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.5cm, row sep=.1ex] DT & NNS & IN & DT & NN & VBP & NNP & .
The & prints & of & every & vase & aggravate & Nina & .
[…]
Figure 8: Distractor agreement relational noun sample well-formed sentence Stanza dependency parse tree. The nsubj edge in green is the critical edge.
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.5cm, row sep=.1ex] DT & NNS & WDT & VBD & VBG & DT & NN & VB & RB & .
All & libraries & that & were & aggravating & a & legislature & joke & around & .
[…]
Figure 9: Distractor agreement relative clause sample well-formed sentence Stanza dependency parse tree. The nsubj edge in green is the critical edge.
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.5cm, row sep=.1ex] DT & NNS & IN & NN & VBP & VBN & NNP & .
The & plays & about & art & have & alarmed & Mitchell & .
[…]
Figure 10: Distractor agreement relative clause sample that we filter out since the critical word is the auxiliary “have”, which is not connected to the subject “plays”.
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.5cm, row sep=.1ex] NNP & VBD & VBN & WP & DT & NN & VBD & .
Marcus & had & remembered & who & some & lady & disliked & .
[…]
Figure 11: Wh. vs. that with gap sample well-formed sentence Stanza dependency parse tree. The obj edge in green is the critical edge.
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.2cm, row sep=.1ex] NNP & VBD & WP & DT & NNS & WDT & VBD & JJ & MD & VB & RP & .
Brenda & knew & what & all & bananas & that & went & bad & can & look & like & .
[…]
Figure 12: Wh. vs. that with gap long distance sample well-formed sentence Stanza dependency parse tree. The obj edge in green is the critical edge.
D.2 Filler-gap
We establish the critical edge for wh. vs. that with gap and wh. vs. that with gap long distance in the same way. Fig 11 and Fig 12 show sample acceptable parse trees for wh. vs. that with gap and wh. vs. that with gap long distance. The critical edge is the edge between the wh-word and the verb it is the direct (obj) or oblique object (obl) of; any other case is filtered out. After filtering, there are 972 valid minimal pairs for wh. vs. that with gap and 885 for wh. vs. that with gap long distance.
Appendix E Probe Training and Comparison
Figure 13: Penn Treebank test set head word probe UAS for each layer of a sample of our models. We train a probe on each layer of the model using the train and validation splits and select the probe for the layer with the best test set accuracy – indicated in the plot with a star icon – for BLiMP evaluation. For most models, the best layer is in the first half of the model.
Figure 14: Correlating $g^{\text{head}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ BLiMP attachment scores. Across all models, $g^{\text{head}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ syntax probes show a significant, moderately positive correlation in accuracy averaged across BLiMP test suites. Despite having an additional directional aspect, $g^{\text{head}}_{\phi}$ trains similarly to $g^{\text{struct}}_{\phi}$.
Figure 15: Correlating $g^{\text{ortho}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ BLiMP attachment scores. Across all models, $g^{\text{ortho}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ syntax probes show a strongly positive correlation in accuracy averaged across BLiMP test suites. Despite the larger capacity of the orthogonal probe, performance largely mirrors its simpler variant, suggesting that the extra dimensions are not critical to learning the parse tree probing task.
Appendix F Full BLiMP and per-Phenomenon OLS Regressions
Dataset Simple Regression ($x_{1}=g^{\text{struct}}_{\phi}$) Multiple Regression ($x_{1},x_{2}=g^{\text{ctrl}}_{\phi}$) LRT
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
Overall 0.133 0.117 0.049 0.109 0.214 0.051 1.161 0.281
Anaphor Agreement 0.165 0.019 0.268 0.168 0.040 0.243 0.027 1.0
Binding 0.398 0.106 0.178 0.386 0.155 0.155 0.18 1.0
Ellipsis -0.421 0.106 0.181 -0.426 0.152 0.153 0.011 1.0
NPI Licensing 0.408 0.121 0.165 0.404 0.155 0.143 0.266 1.0
S-Selection -0.249 0.242 0.125 -0.213 0.374 0.255 6.254 0.149
Island Effects 0.483 0.45 0.087 0.435 0.862 0.065 0.326 1.0
Filler Gap Dependency -0.028 1.0 -0.033 0.033 1.0 -0.05 0.546 1.0
Quantifiers 0.267 1.0 0.008 0.253 1.0 -0.019 0.232 1.0
Determiner Noun Agr -0.063 1.0 -0.013 -0.062 1.0 -0.046 0.067 1.0
Argument Structure 0.198 1.0 0.006 0.164 1.0 0.043 2.295 1.0
Control/Raising 0.118 1.0 0.003 0.101 1.0 0.009 1.295 1.0
Subject Verb Agr -0.108 1.0 0.007 -0.047 1.0 0.128 5.236 0.243
Irregular Forms -0.046 1.0 -0.029 -0.027 1.0 0.277 12.376 0.006
Table 5: Comparison of simple and multiple linear regressions with $x_{1}=$ $g^{\text{struct}}_{\phi}$ UUAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$, extending Table 1. Per-phenomenon rows are sorted in order of increasing simple regression corrected $p$-value. Anaphor agreement is the only one with statistical significance even after incorporating the control. The other phenomena show substantial variation in the strength and direction of association. In particular, we note that the adjusted $R^{2}$ jumps for s-selection and irregular forms when adding the control. The significant $p$-value of the irregular forms LRT is concurrent with Fig 5 showing the control probe to be a moderate predictor of irregular forms minimal pairs.
Dataset Simple Regression ($x_{1}=g^{\text{ortho}}_{\phi}$) Multiple Regression ($x_{1},x_{2}=g^{\text{ctrl}}_{\phi}$) LRT
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
Overall 0.127 0.157 0.035 0.102 0.266 0.041 1.299 0.254
Anaphor Agreement 0.173 0.002 0.364 0.173 0.005 0.342 0.004 1.0
Binding 0.431 0.08 0.195 0.419 0.12 0.173 0.222 1.0
Ellipsis -0.381 0.297 0.124 -0.376 0.414 0.095 0.01 1.0
NPI Licensing 0.326 0.618 0.082 0.32 0.715 0.057 0.215 1.0
S-Selection -0.216 0.727 0.068 -0.175 1.0 0.205 6.164 0.156
Filler Gap Dependency -0.1 1.0 -0.027 -0.058 1.0 -0.049 0.402 1.0
Quantifiers 0.297 1.0 0.015 0.266 1.0 -0.012 0.21 1.0
Determiner Noun Agr -0.089 1.0 -0.004 -0.089 1.0 -0.036 0.077 1.0
Argument Structure 0.174 1.0 -0.003 0.132 1.0 0.032 2.24 1.0
Control/Raising 0.165 1.0 0.041 0.153 1.0 0.048 1.304 1.0
Subject Verb Agr -0.119 1.0 0.014 -0.059 1.0 0.132 5.183 0.251
Island Effects 0.42 1.0 0.045 0.356 1.0 0.028 0.493 1.0
Irregular Forms -0.093 1.0 -0.016 -0.078 1.0 0.287 12.45 0.005
Table 6: Comparison of simple and multiple linear regressions with $x_{1}=$ $g^{\text{ortho}}_{\phi}$ UUAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$, extending Table 2. Per-phenomenon rows are sorted in order of increasing simple regression corrected $p$-value.
Dataset Simple Regression ($x_{1}=g^{\text{head}}_{\phi}$) Multiple Regression ($x_{1},x_{2}=g^{\text{ctrl}}_{\phi}$) LRT
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
Overall 0.0 0.998 -0.033 -0.048 0.700 0.091 5.194 0.023
Binding 0.534 0.064 0.209 0.522 0.108 0.184 0.085 1.0
S-Selection -0.2 1.0 -0.007 -0.255 1.0 0.068 3.553 0.654
NPI Licensing 0.25 1.0 0.006 0.233 1.0 -0.001 0.882 1.0
Filler Gap Dependency -0.125 1.0 -0.028 -0.15 1.0 -0.055 0.256 1.0
Quantifiers 0.434 1.0 0.045 0.385 1.0 0.1 2.98 0.843
Determiner Noun Agr -0.096 1.0 -0.013 -0.098 1.0 -0.046 0.037 1.0
Ellipsis -0.109 1.0 -0.02 -0.1 1.0 -0.053 0.087 1.0
Argument Structure 0.025 1.0 -0.033 -0.037 1.0 0.076 4.658 0.402
Control/Raising 0.241 1.0 0.047 0.201 1.0 0.076 2.083 1.0
Subject Verb Agr -0.205 1.0 0.041 -0.221 1.0 0.025 0.56 1.0
Island Effects 0.392 1.0 -0.001 0.397 1.0 -0.035 0.025 1.0
Anaphor Agreement 0.102 1.0 0.005 0.096 1.0 -0.027 0.069 1.0
Irregular Forms -0.120 1.0 -0.026 -0.099 1.0 0.069 4.217 0.48
Table 7: Comparison of simple and multiple linear regressions with $x_{1}=$ $g^{\text{head}}_{\phi}$ UAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$, extending Table 3. Per-phenomenon rows are sorted in order of increasing simple regression corrected $p$-value. Having 12/13 phenomena with $p$-value 1.0 is a striking result; moreover, comparing with $g^{\text{struct}}_{\phi}$ regressions in Table 5, $g^{\text{head}}_{\phi}$ shows relatively much weaker ability to fit anything at all.
Dataset Simple Regression ($x_{1}=g^{\text{ctrl}}_{\phi}$)
$\beta_{1}$ $p$-value Adj. $R^{2}$
Overall 0.127 0.165 0.032
Irregular Forms 0.287 0.009 0.299
S-Selection -0.445 0.141 0.167
Subject Verb Agr -0.343 0.178 0.151
Binding 0.118 1.0 -0.009
NPI Licensing 0.086 1.0 -0.022
Filler Gap Dependency -0.193 1.0 -0.016
Quantifiers 0.232 1.0 -0.015
Determiner Noun Agr -0.036 1.0 -0.031
Ellipsis -0.134 1.0 -0.018
Argument Structure 0.208 1.0 0.048
Control/Raising 0.161 1.0 0.016
Island Effects 0.202 1.0 0.01
Anaphor Agreement 0.025 1.0 0.005
Table 8: Simple linear regression results for the structural probe control (trained on the best layers for $g^{\text{struct}}_{\phi}$) $x_{1}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon rows are sorted in order of increasing corrected $p$-value. Irregular forms has statistical significance and substantially higher $R^{2}$ than most of the syntax probe simple regressions for any phenomena (Table 5 and Table 7). This critical result shows that even a non-syntactic control can achieve a significant positive association with one syntactic benchmark, which thus urges caution when interpreting the statistical signifiance of anaphor agreement for $g^{\text{struct}}_{\phi}$.
Dataset Simple Regression ($x_{1}=g^{\text{ctrl}}_{\phi}$)
$\beta_{1}$ $p$-value Adj. $R^{2}$
Overall 0.043 0.031 0.117
Argument Structure 0.085 0.5 0.106
Irregular Forms 0.078 0.568 0.096
Quantifiers 0.143 0.865 0.069
S-Selection 0.043 1.0 0.057
Binding 0.038 1.0 -0.007
NPI Licensing 0.038 1.0 -0.001
Filler Gap Dependency 0.027 1.0 -0.027
Determiner Noun Agr 0.002 1.0 -0.033
Ellipsis -0.023 1.0 -0.028
Control/Raising 0.049 1.0 0.053
Subject Verb Agr 0.008 1.0 -0.026
Island Effects -0.004 1.0 -0.033
Anaphor Agreement 0.004 1.0 -0.025
Table 9: Simple linear regression results for the head word probe control (trained on the best layers for $g^{\text{head}}_{\phi}$) $x_{1}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon rows are sorted in order of increasing simple regression corrected $p$-value. While all phenomena have extremely low or near-zero explanatory power that does not achieve statistical significance, the fit on the full dataset is significant but has near-zero effect size.
Appendix G Subject-Verb Agreement & Filler-Gap Experiments
Figure 16: Analogous to Fig 6 but for the head word probe. Conclusion is similar.